OpenStudy (anonymous):

simplify √27x^5 √3x^3

OpenStudy (anonymous):

what is a common number found in each discriminent? we have a 27x^5 and a 3x^3, what can you factor out?

OpenStudy (jdoe0001):

\(\bf \large \sqrt{27}x^5\times \sqrt{3}x^3\quad ?\)

OpenStudy (anonymous):

yes how do you simplify that

OpenStudy (jdoe0001):

\(\bf \sqrt{27}x^5\times \sqrt{3}x^3\implies \sqrt{27\times 3}x^5x^3\) and as you recall from the rules of exponents, same base, you add the exponents, thus \(\bf \sqrt{27}x^5\times \sqrt{3}x^3\implies \sqrt{27\times 3}x^5x^3\implies \sqrt{27\times 3}x^{5+3}\)

OpenStudy (anonymous):

would it be x√3x 3times2x√3x?

OpenStudy (anonymous):

\[x \sqrt{3x}\] \[3\times2x \sqrt{3x}\]

OpenStudy (jdoe0001):

hmmm \(\bf \large \sqrt{27}x^5\times \sqrt{3}x^3\) has the "x" variables outside the radical

OpenStudy (anonymous):

the x^5 and x^3 are both in the square root

OpenStudy (anonymous):

is it like this?